OpenStudy (anonymous):

i have a math problem that is for estimating 6 3/4 + 7 3/4 which i get it estimated to 7 + 8 which gives me 15 but the back of the book says its 14 1/2 can you help me with this problem

OpenStudy (anonymous):

6 + 7 = 13 3/4 + 3/4 = 6/4 = 1 1/2 13 + 1 1/2 = 14 1/2

OpenStudy (anonymous):

Thanks